OpenStudy (thephysicsman):

Bob is planning a picnic for his family of 16 people. His total budget is 900 dollars. in addition to food, he wants to rent a house costing 390 dollars.

OpenStudy (thephysicsman):

which inequality represents how much he can spend per person for food and beverages?

OpenStudy (thephysicsman):

I was thinking that his total expenses can't exceed 900 dollars. so \[$900 \le 390+16x \]

OpenStudy (shamim):

So his total expenditure = 390+16x Right?

OpenStudy (shamim):

He has 900 dollar to spend

OpenStudy (thephysicsman):

yeah he has 900 dollars to spend

OpenStudy (thephysicsman):

the cost of the house is 390

OpenStudy (shamim):

He can not spend more than 900 dollar

OpenStudy (thephysicsman):

yeah

OpenStudy (rsmith6559):

Wouldn't $900 be greater than or equal to (390 + 16x) ?

OpenStudy (shamim):

In ur inequality spending is greater than 900dollar. Right

OpenStudy (thephysicsman):

I think I did something wrong in the inequality

OpenStudy (shamim):

Yes

OpenStudy (thephysicsman):

Is that the proper way to write it if 900 dollars is the maximum he can spend?

OpenStudy (shamim):

Yes

OpenStudy (thephysicsman):

What would this represent? \[900 \ge 390+16x \]

OpenStudy (shamim):

390+16x will b less than or equal to 900 dollar Never 390+16x is greater than 900 dollar

OpenStudy (thephysicsman):

is there any difference usually inequalities are written like x < 3 like in that form not the other way around

OpenStudy (shamim):

Ur inequality represents 390+16x is less than or equal to 900 dollar Right?

OpenStudy (shamim):

Ok wt does it mean by x>6

OpenStudy (shamim):

It means the value of x is greater than 6 Right?!

OpenStudy (shamim):

Can u tell me the meaning of x>4

OpenStudy (thephysicsman):

@shamim that means that x is greater than 4